### v3

# Cross-Subject Data Splitting for Brain-to-Text Decoding

Abstract
Recent major milestones have successfully decoded non-invasive brain signals (e.g. functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG)) into natural language. Despite the progress in model design, how to split the datasets for training, validating, and testing still remains a matter of debate. Most of the prior researches applied subject-specific data splitting, where the decoding model is trained and evaluated per subject. Such splitting method poses challenges to the utilization efficiency of dataset as well as the generalization of models. In this study, we propose a cross-subject data splitting criterion for brain-to-text decoding on various types of cognitive dataset (fMRI, EEG), aiming to maximize dataset utilization and improve model generalization. We undertake a comprehensive analysis on existing cross-subject data splitting strategies and prove that all these methods suffer from data leakage, namely the leakage of test data to training set, which significantly leads to overfitting and overestimation of decoding models. The proposed cross-subject splitting method successfully addresses the data leakage problem and we re-evaluate some SOTA brain-to-text decoding models as baselines for further research.
1 Introduction
Brain-to-text decoding aims to recover natural language from brain signals stimulated by corresponding speech. Recent studies (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023; Duan et al., 2024) have successfully decoded non-invasive brain signals (e.g. fMRI, EEG) to text by applying deep computational neural networks. However, no consensus has reached on how to split the cognitive dataset for training, validating, and testing. Most of the prior work (Ye et al., 2023; Tang et al., 2023) performed subject-specific data splitting for training and evaluating decoding models. Under this splitting rule, data for training, validating, and testing all comes from one specific subject from the cognitive dataset. For example, Tang et al. (2023) picked three subjects out of seven and conducted model training and evaluation on the three subjects respectively. This subject-specific splitting method causes two main problems. First, it only utilizes a tiny part of the whole dataset. Since the collection of brain signals is costly and time-consuming, such splitting method results in significant waste of data resources. Second, it leads to the poor generalization of decoding models. As every subject’s brain has unique functional and anatomical structures, subject-specific models may exhibit considerable variability across individuals and fail to generalize to other subjects. Moreover, decoding models trained from scratch on limited data are prone to facing the overfitting problem.
Figure 1: Illustration of naturalistic language comprehension dataset for brain-to-text decoding. Path with the same color indicates one sample for training/validating/testing.
Some studies (Wang and Ji, 2022; Xi et al., 2023) began to shed light on cross-subject data splitting, which views all the subjects’ data as a whole and performs splitting according to a given ratio (e.g. 8:1:1 for the training set, validation set, and test set). Cross-subject data splitting effectively compensates for the shortcomings of subject-specific splitting and has been widely applied in brain-to-image decoding (Wang et al., 2024; Liu et al., 2024). However, unlike datasets for brain-to-image decoding, where subjects are guided to see different and unrepeated pictures, different subjects will hear the same story in the naturalistic language comprehension dataset for brain-to-text decoding, which challenges cross-subject data splitting. As shown in Figure 1, such dataset is usually formatted in subject-task-text-signal ($S$-$M$-$T$-$F$) pair, indicating the brain signal $F$ of subject $S$ stimulated by hearing text $T$ from task $M$. Current cross-subject data splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized as five categories: (1) split by subjects $S$, (2) split by tasks $M$, (3) split by randomly picking signal frames $F$, (4) split by randomly picking signal frames under certain task $F$-$M$, (5) split by randomly picking consecutive signal frames under certain task $F$-$M$. However, based on our observations, all these splitting methods suffer from data leakage problem, namely part of the test data is mixed into the training set, which leads to overfitting in model training and overestimation in model evaluation.
Specifically, modern brain-to-text decoding models follow an encoder-decoder manner. We pick two representative models: EEG2Text (Wang and Ji, 2022) and UniCoRN (Xi et al., 2023) for investigating the damage of data leakage. The former is an end-to-end encoder-decoder framework, while the latter first pre-trains the encoder and then applies it in the decoder training. Experiments support that data leakage affects model training on both the encoder side and decoder side. For the encoder component, if subjects’ brain signals in the test set are mixed into the training set, the encoder will become overfitting and fail to well represent unseen subjects’ brain signals. As to the decoder, the situation gets worse if text stimuli are leaked. Since the decoder follows an auto-regressive manner and generates token by token, data leakage will cause the decoder to memorize seen paragraphs during the teacher-forcing training stage, which leads to poor generalization to unseen text.
To avoid data leakage and fairly evaluate the performance of brain-to-text decoding models, we propose a cross-subject data splitting criterion. We focus on fMRI and EEG signals in this study, although the proposed criterion could be applied to any datasets satisfying the prescribed format. In the proposed method, the dataset is split according to subject-text ($S$-$T$) pairs with the following rules: (1) Brain signals collected from specific subject in validation set and test set will not appear in training set, which means the trained encoder cannot get access to any brain information belonging to subjects in test set. (2) Text stimuli in validation set and test set will not appear in training set. The decoder learns to reconstruct language with brain signals instead of memorizing seen text.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we propose the first cross-subject data splitting criterion for brain-to-text decoding.
  • We comprehensively analyze current cross-subject data splitting methods and find that all existing methods suffer from data leakage problem, which severely affects the training and evaluation of decoding models.
  • Some SOTA brain-to-text decoding models are re-evaluated under the proposed cross-subject data splitting method as baselines for further research.
Figure 2: Different splitting methods for cognitive dataset. (Color printing is preferred.)
2 Related Work
Brain Signal
Brain signals can be classified into three categories: invasive, partially invasive, and non-invasive according to how close electrodes get to brain tissue. In this paper, we mainly focus on non-invasive signals EEG and fMRI. EEG signal is electrogram of the spontaneous electrical activity of the brain, with frequencies ranging from 1 Hz to 30 Hz. EEG is of high temporal resolution and relatively tolerant of subject movement, but its spatial resolution is low and it can’t display active areas of the brain directly. fMRI measures brain activity by detecting changes of blood flow. Blood flow of a specific region increases when this brain area is in use. The spatial resolution of fMRI is measured by the size of voxel, which is a three-dimensional rectangular cuboid ranging from 3mm to 5mm (Vouloumanos et al., 2001; Noppeney and Price, 2004). Unlike EEG which samples brain signals continuously, fMRI samples based on a fixed time interval named TR, usually at second level.
Brain-to-text Decoding
Previous research on brain-to-text decoding (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021). These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between brain signals and words in vocabulary. Despite relatively good accuracy, these methods fail to generalize to unseen words. Some progress (Sun et al., 2019) has been made by expanding word-level decoding to sentence-level through encoder-decoder framework or using less noisy ECoG data (Burle et al., 2015; Anumanchipalli et al., 2019). However, these models struggle to generate accurate and fluent sentences limited by decoder ability. Wang and Ji (2022) introduced the first open vocabulary EEG-to-text decoding model by leveraging the power of pre-trained language models. Xi et al. (2023) improved the model design and proposed a unified framework for decoding both fMRI and EEG signals.
3 Methodology
In this section, we will first introduce the formal definition of brain-to-text decoding and the general description of dataset format. Then we systematically analyze current cross-subject data splitting methods and point out that all existing methods suffer from two kinds of data leakage issues: brain signal leakage and text stimuli leakage. Finally, a cross-subject splitting criterion is proposed which avoids the above-mentioned data leakage problems.
3.1 Task Definition
Given the brain signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$. The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG. The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously. To fMRI, consistent text segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$, where $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$, and $|T_{j}|=|F_{ij}|=L$. To EEG, since signals corresponding to complete text stimuli are available and they are continuous, we bond text $T_{j}$ (i.e. text stimuli) and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Under most scenarios, each text stimulus $T_{j}$ belongs to one certain task $M_{k}$. So the signal-text pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
Figure 3: The process of our proposed cross-subject data splitting method. (Color printing is preferred.)
One of the purposes of cross-subject splitting is to endow models the ability to decode unseen subject’s brain signal. As a result, if brain signal $F_{ij}$ appears in test set $S_{test}$, any signal $F_{i*}$ belonging to subject $i$ should not appear in training set $S_{train}$. Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$. The above splitting rules for training set can be formulated by Cartesian product:
$S_{train}=F_{train}\times T_{train},$ (1)
where
$F_{train}=\{F_{ij}|i\in I\},$ (2)
$I=\{i|F_{ij}\notin S_{val},S_{test},\forall j\},$ (3)
and
$T_{train}=\{T_{kj}|T_{kj}\notin S_{val},S_{test}\}.$ (4)
Such rules are also applicable to validation and test set. We omit their displays here for simplicity.
3.2 Current Splitting Methods
Current cross-subject data splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$. More specifically, the five dataset splitting methods are characterised as (1) split by subjects $S_{i}$, (2) split by tasks $M_{k}$, (3) split by randomly picking signal frames $F_{ij}$, (4) split by randomly picking signal frames under certain task $F_{ij}$-$M_{k}$, (5) split by randomly picking consecutive signal frames under certain task $F_{ij}$-$M_{k}$, corresponding to (a), (b), (c), (d), (e) in Figure 2. Figure 2 vividly displays the differences between current dataset splitting methods. In this example, we choose 4 subjects ($S_{1}$ to $S_{4}$) with 3 tasks ($M_{1}$ to $M_{3}$) each containing 4 ($T_{11}$ to $T_{14}$), 3 ($T_{21}$ to $T_{23}$), 4 ($T_{31}$ to $T_{34}$) text stimuli respectively. $F_{ij}$ indicates the brain signal of $i$-th subject stimulated by $j$-th text under task $M_{k}$, e.g. $F_{21}$ means brain signal of $S_{2}$ hearing $T_{11}$.
The line connecting two symbols indicates they are related to one sample in dataset. Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it indicates that subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$. Some symbols are connected with several lines. For example, the four lines between $S_{1}$ and $M_{1}$ correspond to $\left\langle M_{1},T_{11}\right\rangle$, $\left\langle M_{1},T_{12}\right\rangle$, $\left\langle M_{1},T_{13}\right\rangle$, $\left\langle M_{1},T_{14}\right\rangle$ counting from left to right. Similarly, the three lines between $M_{1}$ and $T_{11}$ correspond to $\left\langle S_{1},M_{1}\right\rangle$, $\left\langle S_{2},M_{1}\right\rangle$, $\left\langle S_{3},M_{1}\right\rangle$ respectively. The same rules can be extended to other lines and symbols. The green lines and orange lines stand for training samples and testing samples. The grey dotted line means the sample is abandoned, which will be introduced in our data splitting method. As the splitting of validation set is the same as test set, we only consider training set and test set in this section for simplicity.
We will use method (a), (b), (c), (d), (e) to represent five current dataset splitting methods in the rest of the paper. Method (a) splits the dataset according to subjects, which can be described as
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|S_{i}\notin S_{val},S_{test}\}$ (5)
for training set. Method (b) splits the dataset according to tasks, which is described as
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|M_{k}\notin S_{val},S_{test}\}$ (6)
for training set. Method (c), (d), and (e) all split the dataset according to brain signal frames
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}.$ (7)
However, there are slight differences between these three methods. Method (c) views all the brain signal frames in dataset as a whole and splits according to the default proportion (e.g. 8:1:1). Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then unions all training sets under different tasks to form a complete set for training. Method (e) is similar to method (d). They both first split training, validation, and test set under certain task proportionally and then union them. The difference lies in that method (d) randomly picks signal frames, while method (e) picks consecutive signal frames.
Type Method Narratives / ZuCo Average
seed1111 seed2222 seed3333 seed4444
[…]
(b) 6.73 / - 6.32 / - 7.7 / - 17.93 / - 9.67 / -
(c) 12.55 / 12.52 12.52 / 12.55 12.48 / 12.48 12.44 /12.46 12.50 / 12.50
(d) 12.81 / 12.60 12.8 / 12.58 12.78 / 12.56 12.79 / 12.61 12.795 / 12.59
(e) 12.28 / - 12.27 / - 12.26 / - 12.27 / - 12.27 / -
(f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
[…]
(b) 0.00 / - 0.00 / - 0.00 / - 0.00 / - 0.00 / -
(c) 100.00 / 13.21 100.00 / 13.06 100.00 / 12.91 100.00 / 13.1 100.00 / 13.07
(d) 99.93 / 0.00 99.81 / 0.00 99.54 / 0.00 99.99 / 0.00 99.82 / 0.00
(e) 9.19 / - 9.31 / - 9.36 / - 9.29 / - 9.29 / -
(f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
Table 1: Results of Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR).
We first point out the data leakage problems in current splitting methods through the analysis of training set and test set composition. Specifically, following the definition in subsection 3.1, two kinds of data leakage, brain signal leakage and text stimuli leakage, are defined. The data leakage situation of different methods can be reflected through visualization in Figure 2. Lines between $S_{i}$ and $M_{k}$ indicate brain signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation. If lines associated with $S_{i}$ or $T_{kj}$ are of different colours, data in test set leaks into training set. Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two kinds of brain signal too. Since fMRI signals need to be sampled continuously with a certain length $L$, one sample shown in Figure 2 (e.g. $S_{1}$-$M_{1}$-$T_{11}$-$F_{11}$) is actually the first part of one fMRI sample, with $L-1$ continuous brain signal frames following (e.g. $F_{12},F_{13},\ldots$). In this sense, brain signal leakage doesn’t exist in method (a) for EEG, but method (a) suffers from text stimuli leakage. Text stimuli leakage does not exist in method (b) but brain signal leakage exists. Method (c) suffers from both brain signal leakage and text stimuli leakage. Method (d) and (e) are the same to EEG, with brain signal leakage. For fMRI, the situation of data leakage for different methods is similar to EEG, except for method (d) and (e), which are the same for EEG but actually different for fMRI. Method (d) suffers from both brain signal leakage and text stimuli leakage while in method (e) text stimuli leakage happens in the overlap between training samples and test samples.
3.3 Cross-Subject Splitting Criterion
To eliminate data leakage from both brain signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{j}\right\rangle$ pairs as shown in (f) of Figure 2. Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two data splitting methods. As to EEG dataset where $F_{ij}$ and $T_{j}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$. $\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},T_{j}\right\rangle$ pair in the dataset. $N$ is the total number of subjects and $M$ is the total number of unique text stimuli. We assert $M>N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject. As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$. Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$. In step 4, some edges satisfying zero data leakage condition are not included in the graph. We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
The sample in fMRI dataset is formatted in $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$ and $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$. If we follow the same process as EEG, text stimuli leakage will occur in the overlapping part of two samples, when one sample is assigned to training set and the other is assigned to validation or test set. We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data leakage. Instead of $\left\langle S_{i},T_{j}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the above-mentioned algorithm. More details and pseudo-code are available in Appendix B.
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
UniCoRN 10+1e-3+(a) 49.56 30.49 21.07 15.49 44.83 50.41 40.65
10+1e-3+(b) 26.37 7.50 2.48 0.99 22.28 25.99 19.62
10+1e-3+(c) 50.24 30.83 21.23 15.60 44.68 49.44 41.01
10+1e-3+(d) 49.63 30.29 20.85 15.32 45.06 50.47 41.03
10+1e-3+(e) 28.94 9.39 4.07 1.53 21.68 24.64 19.49
UniCoRN 20+1e-4+(a) 50.19 34.25 25.98 21.00 46.59 50.36 43.62
30+1e-4+(a) 55.46 40.99 32.85 27.56 52.08 55.02 49.68
20+1e-4+(b) 25.91 8.80 3.84 1.66 20.65 27.74 16.57
30+1e-4+(b) 25.91 8.80 3.84 1.66 20.65 27.74 16.57
20+1e-4+(c) 72.44 60.84 53.35 47.88 70.52 74.10 67.53
30+1e-4+(c) 72.82 61.42 53.95 48.44 71.24 74.41 68.57
20+1e-4+(d) 65.31 51.02 42.54 36.72 62.76 67.09 59.29
30+1e-4+(d) 66.56 53.00 44.75 39.02 63.89 67.51 60.95
20+1e-4+(e) 32.15 12.34 5.57 2.45 24.28 30.43 20.35
30+1e-4+(e) 32.15 12.34 5.57 2.45 24.28 30.43 20.35
Table 2: Generation quality of UniCoRN model for fMRI under different training settings. Here UniCoRN indicates the encoder of UniCoRN is randomly initialized instead of pre-trained through signal reconstruction task.
4 Experimental Settings
We test some SOTA brain-to-text decoding models on two popular cognitive datasets Narratives (Nastase et al., 2021) and ZuCo (Hollenstein et al., 2018). Dataset details are introduced in Appendix A. Comprehensive experiments are conducted to evidence the existence of the following phenomena: (1) Brain signals and text stimuli in test set leak into training set in all current dataset splitting methods. (2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data leakage. Because the number of tasks in EEG dataset is too small and method (e) makes no difference to EEG as method (d), we only consider method (a), (c), (d) for EEG.
4.1 Implementation
We follow the same settings of UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022), except all the datasets are split to the ratio of 8:1:1 for fair comparison. All experiments are conducted on NVIDIA A100-SXM4-40GB GPUs. More details are shown in Appendix A.
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
UniCoRN 50+1e-4+(a) 58.09 49.23 43.23 38.43 63.88 61.12 67.50
80+1e-4+(a) 60.88 50.52 43.42 37.84 65.17 61.16 70.72
50+1e-4+(c) 52.30 42.89 36.80 32.17 57.39 51.09 67.29
80+1e-4+(c) 60.78 55.92 53.18 51.10 84.64 63.16 71.50
50+1e-4+(d) 22.90 7.36 2.71 0.95 17.73 19.90 17.33
80+1e-4+(d) 22.90 7.36 2.71 0.95 17.73 19.90 17.33
EEG2Text 50+1e-4+(a) 51.22 33.83 22.99 16.05 46.40 46.85 46.58
80+1e-4+(a) 63.32 52.52 45.19 39.50 65.96 64.74 68.01
50+1e-4+(c) 53.83 38.99 29.57 23.01 53.64 54.19 53.56
80+1e-4+(c) 65.42 57.56 52.56 48.60 73.00 69.99 77.01
50+1e-4+(d) 23.92 8.16 3.21 1.20 20.78 19.96 23.89
80+1e-4+(d) 23.92 8.16 3.21 1.20 20.78 19.96 23.89
Table 3: Generation quality of UniCoRN and EEG2Text model for EEG under different training settings.
4.2 Data Leakage Metrics
We have analyzed two kinds of data leakage: brain signal leakage and text stimuli leakage in section 3. In this part, we will quantify two kinds of data leakage through experiments.
To better illustrate the extent of data leakage of different data splitting methods, we design two novel evaluation metrics named Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR) for detecting brain signal leakage and text stimuli leakage. Note that the situation for validation set is the same as test set, so we only consider test set in experiments. BSLR indicates the average percentage of each subject’s brain signals in test set appearing in training set, which could be formulated as
$\frac{1}{N}\sum_{i=1}^{N}\min(1,\frac{|\{F_{ij}|F_{ij}\in S_{test}\cap S_{train}\}|}{|\{F_{ij}|F_{ij}\in S_{train}\}|})$ (8)
where $N$ stands for the total number of subjects in test set. ||$\cdot$|| ⋅ | stands for the cardinality of a set. Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than one.
The definition of TSLR is somewhat different for EEG signal and fMRI signal. As to EEG signal where brain signals are sampled continuously, it’s easy to match certain text stimuli with corresponding signals. Its TSLR is similar to BSLR, which indicates the average percentage of certain text in test set appearing in training set. TSLR for EEG data can be calculated through
$\frac{1}{M}\sum_{j=1}^{M}\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{train}\}|})$ (9)
where $M$ stands for the total number of unique text periods in test set and $T_{ij}$ stands for $j$-th period of text stimuli received by $i$-th subject.
The fMRI signal is sampled discretely with a deterministic interval TR, making it hard to acquire signals corresponding to text. Previous methods instead concatenated continuous fMRI frames of certain length with their corresponding text segments as training samples. As a result, we consider the average percentage of the same text segments in test set appearing in training set as TSLR for fMRI signal. It can be formulated as
$\frac{1}{M}\sum_{j=1}^{M}\tau\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|S_{test}|\times L}$ (10)
where $\tau=0$ if $\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}=\emptyset$ else
$\tau=\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}).$ (11)
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
Narratives UniCoRN 22.83 5.69 1.43 0.48 15.55 24.80 19.04
ZuCo UniCoRN 23.32 7.78 3.01 1.09 18.47 20.00 17.92
EEG2Text 24.49 7.49 2.28 0.62 23.98 23.95 25.74
Table 4: A fair benchmark for evaluating brain-to-text decoding.
5 Results and Analysis
5.1 Verification for Data Leakage
We test current data splitting methods and our method on fMRI dataset “Narratives” and EEG dataset ZuCo. Considering the influence of randomness in splitting, we randomly select four seeds for experiments. The results are shown in Table 1, and are consistent with theoretical analysis. For fMRI, current methods apart from method (a) suffer from brain signal leakage, while method (a) has serious text stimuli leakage. Method (b) gets no text stimuli leakage but has slight brain signal leakage. The situation for EEG is similar to that of fMRI. Apart from our proposed method (f), there is no way to achieve zero brain signal leakage and text stimuli leakage at the same time.
5.2 Damage of Data Leakage
Brain signal leakage and text stimuli leakage will damage brain-to-text decoding models from both the encoder side and decoder side.
Effect on Encoder
The encoder of current brain-to-text decoding models can be trained in two ways: either jointly with the decoder or solely pre-trained through a reconstruction task. In the former end-to-end training scenario, it is hard to evaluate encoder performance separately. So we mainly focus on the latter, in which case the encoder is pre-trained through an encoder-decoder framework to reconstruct input brain signals. The decoder here does not refer to the decoder for text generation. It is similar to the structure of the encoder and will be abandoned once the encoder is pre-trained. Since a proper evaluation index of the encoder’s representation ability is missing, validation loss is applied to measure the effect of data leakage.
We test different splitting methods on two cognitive datasets. The validation loss of encoder is shown in Figure 4. For fMRI, influenced by leakage of brain signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs. The encoder is actually overfitting and degrading. For method (a) and (f) without brain signal leakage, the validation loss quickly rises after reaching the lowest point with a few epochs, satisfying the basic rule of machine learning. For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of brain signal leakage or not. We think the poor spatial resolution of EEG signal might lead to this phenomenon.
Figure 4: Validation loss of encoder under different dataset splitting methods in two datasets.
Effect on Decoder
All SOTA models choose the pre-trained language model BART (Lewis et al., 2020) as decoder. The powerful model is able to achieve fluent open vocabulary text generation. However, if data leakage occurs, due to the auto-regressive generation that calculates the probability of current token based on all previous tokens, the decoder will generate seen text given the first few words, and fail to generalize to unseen text.
The influence of text stimuli leakage on decoder is detected through BLEU scores (Papineni et al., 2002) and ROUGE-1 scores (Lin, 2004), which measure text similarity between generated text and ground truth. If evaluation indicators keep improving as training epochs increase, we believe part of the test set is leaked into training set and the model is overfitting. For fMRI signal, we test five current dataset splitting methods under different training settings. As shown in Table 2, we test two kinds of UniCoRN models. One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is UniCoRN with a randomly initialized encoder. Empirically, the former will perform much better than the latter. However, in method (a), (c), (d), due to text stimuli leakage, if we reduce the learning rate and extend training epochs, UniCoRN performs much better than UniCoRN and its performance keeps rising with longer training epochs. As to method (b) and (e) with no text stimuli leakage, changing training epochs or learning rates makes no obvious difference to model performance. For EEG signal, the conclusion is similar as shown in Table 3. For method (a) and (c) with text stimuli leakage, model performance keeps rising with longer training epochs. For method (d) without text stimuli leakage, both models reach optimal performance after the first few rounds of training epochs.
5.3 A Fair Benchmark
We re-evaluate current SOTA models for brain-to-text decoding under our cross-subject data splitting method and release a fair benchmark. UniCoRN is tested for both fMRI and EEG decoding, EEG2Text model is tested for EEG decoding. The results are listed in Table 4. For EEG dataset, UniCoRN achieves higher results in BLEU-2,3,4 while EEG2Text is better in BLEU-1 and ROUGE-1.
6 Conclusion
In this paper, we conduct a comprehensive study on existing cross-subject data splitting methods, and evidence that all these methods suffer from data leakage problem. Such data leakage largely exaggerates model performance and leads to poor generalization. To fix this issue, we propose a cross-subject data splitting criterion for brain-to-text decoding, aiming to improve the utilization efficiency of cognitive dataset and the generalization ability of decoding models. Experiments are conducted on fMRI and EEG dataset respectively. Current SOTA models are re-evaluated under this proposed splitting method and a fair benchmark is released for further research in this domain.
Limitations
The “Narratives” dataset and the ZuCo dataset provide researchers with precise brain signal resources stimulated by text or voice. However, in the brain-to-text decoding task, both subject’s brain signals and text stimuli in the validation and test set need to be invisible to the training set, which makes splitting these public datasets difficult. Our proposed dataset splitting method meets the above requirements at the expense of discarding some data in the dataset. We recommend future datasets in this domain follow these guidelines. The division of the training set, validation set, and test set should be provided when the dataset is released. Besides, we suggest hiring new subjects with unique stimuli for the validation set and test set, which is good for testing the generalization ability of models without loss of data. What’s more, we find existing models rely more on a strong auto-regressive decoder to achieve good generation quality. The encoder is of limited use in all SOTA models. And we also notice in experiments that the encoder of EEG2Text keeps overfitting whether with or without brain signal leakage. We leave it to research in the future.
Ethics Statement
In this paper, we introduce a new dataset splitting method to avoid data leakage for decoding brain signals to text task. Experiments are conducted on the publicly accessible cognitive datasets “Narratives” and ZuCo1.0 with the authorization from their respective maintainers. Both datasets have been de-identified by dataset providers and used for researches only.
Appendix A Implementation Details
[…]
More details in experiments are supplemented in this section. We perform the same filtering steps to “Narratives” dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022). In BSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively. In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively. The learning rate is set as 1e-4 and 1e-3 separately as the author claimed in the original paper. In the fair benchmark, for fMRI data, encoder of UniCoRN is trained through 1e-4 learning rate and decaying to 1e-6 finally for 30 training epochs. Decoder is trained through 1e-4 learning rate and decaying to 1e-6 finally for 10 training epochs with 90 batch size. Sample length $L$ is set as 10 for all experiments related to fMRI. For EEG data, EEG2Text model is trained with 1e-6 learning rate for 80 epochs. UniCoRN model is trained with the same settings as fMRI data.
Appendix B Details of Cross-Subject Splitting
For fMRI dataset, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{M_{k}\}_{k=1}^{M}$. $\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},M_{k}\right\rangle$ pair in the dataset. $N$ is the total number of subjects and $M$ is the total number of tasks. We assert $M<N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject. As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$. Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$. In step 4, some edges satisfying zero data leakage condition are not included in the graph. We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
We also release the pseudo-code of two dataset splitting methods for EEG and fMRI signal. As shown in Figure 3, our proposed dataset splitting method consists of four steps. The blue lines stand for the situation of original dataset. The main difference between two methods lies in the how $\mathcal{G}_{2}$ is generated. We always choose the side with fewer nodes in bipartite graph $\mathcal{G}_{1}$ to perform $\mathcal{G}_{2}$ generation. For example, in Algorithm 1 where we assert $|\mathcal{U}|<|\mathcal{V}|$, the adjacency matrix is initialized as $M\times N$. In Algorithm 2 where $|\mathcal{V}|<|\mathcal{U}|$, the adjacency matrix is initialized as $N\times K$. All assertions are based on analysis of cognitive datasets.
One more thing to notice is that in Line 14 of both pseudo-code, the loop indicates extending training set, validation set, and test set respectively. So the names of variable should be alternated in the repeat loop and the displayed part in pseudo-cod is a case example of extending training set. We write it in this way for simplicity of expression.
1 Initialize: Bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$, $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$ and $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$, Adjacency matrix $A_{1}$ of $\mathcal{G}_{1}$ where $A_{1}[i][j]=1$ if node $i$ and node $j$ is connected else $A_{1}[i][j]=0$, Adjacency matrix $A_{2}$ of $\mathcal{G}_{2}$ where $A_{2}[i][j]=0$, Array $C$ where $len(C)=len(\mathcal{U})$ and $C[i]=0$;
2 for $u\leftarrow U_{1}$ to $U_{N}$ do
3       $C_{copy}\leftarrow C$;
4       for $v\leftarrow A_{1}[u][0]$ to $A_{1}[u][M]$ do
5             if $v=0$ then
6                  $C_{copy}[v.index]\leftarrow\infty$;
7
8      Minimum $=\min(C_{copy})$;
[…]
11Split by subjects $\mathcal{U}$ according to default ratio;
12 $\mathcal{G}_{2}=\mathcal{G}_{train}\cup\mathcal{G}_{val}\cup\mathcal{G}_{test},\,\mathcal{U}_{train}\cap\mathcal{U}_{val}\cap\mathcal{U}_{test}=\emptyset,\,\mathcal{V}_{train}\cap\mathcal{V}_{val}\cap\mathcal{V}_{test}=\emptyset$;
13 repeat // To three sets respectively, below is for training set
14       for $u$ in $\mathcal{U}$  do
15             for $v$ in $\mathcal{V}$ do
16                   if $e=(u,v)\in\mathcal{E}$ and $e=(u,v)\notin\mathcal{E}^{\prime}_{train}$ and $u\notin\mathcal{U}_{val}\cup\mathcal{U}_{test}$ then
17                        $\mathcal{E}_{train}^{\prime}\leftarrow\mathcal{E}_{train}^{\prime}\cup\{e\}$;
18
19
20
21until $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ are all extended;
22return $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$;
Algorithm 1 Dataset splitting method for EEG signal
23 Initialize: Bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$, $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{M_{k}\}_{k=1}^{K}$, Adjacency matrix $A_{1}$ of $\mathcal{G}_{1}$ where $A_{1}[i][j]=1$ if node $i$ and node $j$ is connected else $A_{1}[i][j]=0$, Adjacency matrix $A_{2}$ of $\mathcal{G}_{2}$ where $A_{2}[i][j]=0$, Array $C$ where $len(C)=len(\mathcal{V})$ and $C[i]=0$;
24 for $v\leftarrow V_{1}$ to $V_{K}$ do
25       $C_{copy}\leftarrow C$;
26       for $u\leftarrow A_{1}[v][0]$ to $A_{1}[v][K]$ do
27             if $u=0$ then
28                  $C_{copy}[u.index]\leftarrow\infty$;
29
30      Minimum $=\min(C_{copy})$;
[…]
33Split by tasks $\mathcal{V}$ according to default ratio;
34 $\mathcal{G}_{2}=\mathcal{G}_{train}\cup\mathcal{G}_{val}\cup\mathcal{G}_{test},\,\mathcal{U}_{train}\cap\mathcal{U}_{val}\cap\mathcal{U}_{test}=\emptyset,\,\mathcal{V}_{train}\cap\mathcal{V}_{val}\cap\mathcal{V}_{test}=\emptyset$;
35 repeat // To three sets respectively, below is for training set
36       for $v$ in $\mathcal{V}$  do
37             for $u$ in $\mathcal{U}$ do
38                   if $e=(u,v)\in\mathcal{E}$ and $e=(u,v)\notin\mathcal{E}^{\prime}_{train}$ and $v\notin\mathcal{V}_{val}\cup\mathcal{V}_{test}$ then
39                        $\mathcal{E}_{train}^{\prime}\leftarrow\mathcal{E}_{train}^{\prime}\cup\{e\}$;
40
41
42
43until $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ are all extended;
44return $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$;
Algorithm 2 Dataset splitting method for fMRI signal